OpenStudy (anonymous):

x^3- 5x^2- 3x+15=0 Solve for x.

hero (hero):

\[x^3 - 5x^2 - 3x + 15 = 0 \\x^2(x - 5) - 3(x - 5) = 0 \\(x - 5)(x^2 - 3) = 0 \\x - 5 = 0 \\x^2 - 3 = 0 \\x = 5 \\x^2 = 3 \\x = ±\sqrt{3}\]

OpenStudy (anonymous):

Thanks!